I need your help with this please.
Solve the following system by substitution. If a unique solution does not exist, state whether the system is inconsistent or dependent. 5x – 2y = -5 and y – 5x = 3

Answer by stanbon(75887)

1st: 5x – 2y = -5
2nd: y – 5x = 3
3rd: 5x-2(5x+3)=-5
5x-10x-6=-5
-5x=1
x=-1/5
Substitute into 2nd to solve for "y", as follows:
y-5(-1/5)=3
y=2
Check in 1st:
5(-1/5) - 2(2)=-5
-1-4=-5
Cheers,
Stan H.